User Tools

Site Tools


lbaops:lbamar2015:v516cholog

At around UT 12:17 Chris Phillips noticed that Ho was stuck on source 11:13:01-35:49:52 - it was TRACKING but never seemed to change targets when the other antennas did.

He reckoned it was stuck at around 10:55 UT.

JMc went out to Mt PLeasant and fixed it at around 13:15 UT. Apparently a sampler died, & it's hung collecting a failed tsys reading.

lbaops/lbamar2015/v516cholog.txt · Last modified: 2015/12/18 16:38 (external edit)